You need to prepare 5.00 L of an aqueous solution Please help
Answer #1

The Ca(OH)2 solution has a pH of 12.50, this means that the pOH is:
pOH = 14-12.5 = 1.5
[OH-] = 10-1.5 = 0.0316 M

Now, that we know this, we know that the solution is dissolved this way:
Ca(OH)2 <--------> Ca2+ + 2OH-

This means that to get the concentration of Ca(OH)2:
[Ca(OH)2] = 0.0316 / 2 = 0.0158 M

the moles and then mass:
moles = 0.0158 mol/L * 5 L = 0.079 moles
mass = 0.079 moles * (40+32+2) = 5.846 g needed to prepare this solution.

The other question post it in another question thread :). Hope this helps